Convert 9.02*10^12 molecules ozone per cubic centimeter into ppb given pressure is 1atm and temperature is 298K.
Answer #1

PV= nRT

P = 1 atm

R = 0.082 lit atm / mol /K

T = 298 K

n = 10^3/ 6.023 *10^23 = 1.66 * 10-21 mole

V = nRT / P....

or. V = 4.06 * 10-20 lit = 1.497 * 10-32billion cube meter[ 1 lit = 10-12 billion cube meter]

hence, 9.02*10^12 molecules in 1.497 * 10-32 billion cube mete

or, 6.025 * 10^44 molecules per billion cube meter